Given that x * y = 4x-2y find the value of (2 * 3) * 4
Found 2 solutions by Alan3354, greenestamps:
Answer by Alan3354(69443)
Answer by greenestamps(13203)


The earlier solution clearly misinterpreted the problem.

Using rules for order of operations, we need to evaluate (2*3) first. Using the given definition of "*",
2*3 = 4(2)-2(3) = 8-6 = 2

Then
2*4 = 4(2)-2(4) = 8-8 = 0